RawlsK29:

If y = 5x − 4, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, −4), (2, 6), (4, 20)} {(0, −4), (2, 6), (4, 16)} {(0, 4), (2, 6), (4, 16)} {(0, 4), (2, −6), (4, 20)}

RawlsK29:

I still don't understand how to do this stuff. @Shadow @Allison @Elsa213

Vocaloid:

take it one point at a time, one answer choice at a time (0, -4) this means that the input of 0 gives an output of -4 to test whether this is true, let x = 0 and see if y = 5x − 4 gives -4 as its output if yes: the point is on the line, test the next point within the same answer choice if not: the point is not on the line, move to the next answer choice

RawlsK29:

Oh okay, so that's not the answer?

Vocaloid:

I haven't said anything about whether it's the answer or not, you still have to go through the calculations see if x = 0 gives an output of -4

RawlsK29:

Okay, that one is not it. Because 5 + -4 = 1 not 0

Vocaloid:

x is the input not the output if x = 0, y = 5(0) - 4 = ?

RawlsK29:

1?

Vocaloid:

5*0 = ?

RawlsK29:

I am so confused right now.

Vocaloid:

what is 5 times 0 ?

RawlsK29:

5

RawlsK29:

i mean 0

Vocaloid:

good, so 5(0) - 4 = -4 meaning that (0,-4) is one of the correct ordered pairs, meaning we can eliminate C and D as possible answers now we test whether (4,20) or (4,16) is an ordered pair on the line

Vocaloid:

y = 5x - 4 if x = 4, what does y equal?

RawlsK29:

umm... i don't know. I feel like an idiot.

Vocaloid:

x = 4 means you can replace "x" with 4" when x shows up so y = 5x - 4 = 5(4) - 4 = ?

RawlsK29:

5(4) that equals 20

Vocaloid:

good, and then subtract 4 from that

RawlsK29:

16

Vocaloid:

awesome, so (4,16) is another point on the graph eliminating A and making B the correct answer

RawlsK29:

(:

RawlsK29:

So, the answer is B? or is there another step?

Vocaloid:

it's B, that's all

RawlsK29:

Thank you, so much! ur a life saver.

RawlsK29:

LOl!